Twice cerilos age, increased by 4 is 22. How old is cerilo after 3 years













Answer by mananth(16946)

Twice cerilos age, increased by 4 is 22. How old is cerilo after 3 years
Twice cerilos age,(c) increased by 4 is 22.
Twice cerilos age,(c) =2c
increased by 4
2c+4
is 22.
2c+4=22
2c +4-4 =22-4
2c=18
/2
c=9
After 3 years cerillo is 9+3 =12 years